OpenStudy (anonymous):

Help Please! How to find the derivative algebraically of f(x)=x^3 at x=-2

OpenStudy (isaiah.feynman):

Definition of the derivative. \[\lim_{x \rightarrow a}\frac{ f(a+h)-f(a) }{ x-a }\]

OpenStudy (isaiah.feynman):

Where a =-2

OpenStudy (anonymous):

So what do I do with the x^3 part?

OpenStudy (isaiah.feynman):

Let x = a+h, but don't substitute a =-2 in this part. Substitute it in f(a)

OpenStudy (isaiah.feynman):

Hey, sorry don't use that formula. Please use this one its more convenient.\[\lim_{x \rightarrow a}\frac{ f(a)-f(x) }{ x-a }\]

OpenStudy (isaiah.feynman):

a=-2. So f(a)=f(-2) and f(x) is x^3.

OpenStudy (anonymous):

I'm sorry. I must still be missing something. I know that the correct answer is 12, but I can't seem to get that answer. I keep getting 0.

OpenStudy (isaiah.feynman):

Did you use the second formula I gave you?

OpenStudy (anonymous):

Yea. It would be (-8)-(x^3)/x^3-(-2), right? I don't know why I'm so confused right now.

OpenStudy (isaiah.feynman):

Oh my God, IT'S ALL MY FAULT!!! What's wrong with me today??!!

OpenStudy (anonymous):

Huh?

OpenStudy (isaiah.feynman):

I wrote the second formula the wrong way. Its supposed to be \[\lim_{x \rightarrow a} \frac{ f(x)-f(a) }{ x-a }\]

OpenStudy (anonymous):

I'm still getting 0 as an answer.

OpenStudy (isaiah.feynman):

So to substitute, \[\lim_{x \rightarrow -2}\frac{ x^{3}-(-8) }{ x-(-2) }\]

OpenStudy (anonymous):

But then you'll get (-8)-(-8) as the numerator and (-2)-(-2) as the denominator, both equaling 0, wouldn't you?

OpenStudy (isaiah.feynman):

No. That expression becomes \[\lim_{x \rightarrow -2}\frac{ x^{3}+8 }{ x+2 }\] Now solve that equation.

OpenStudy (anonymous):

But when you substitute -2 for x, you get (-2)^3, which is -8, and -2. Making it -8+8 over -2+2. Right?

OpenStudy (isaiah.feynman):

You DON'T substitute x for -2!!! That would make the denominator zero and division by zero is undefined. That arrow means x approaches -2.

OpenStudy (anonymous):

So how will the answer end up being 12?

OpenStudy (isaiah.feynman):

I factor the denominator \[\lim_{x \rightarrow -2} \frac{ (x+2)(x^{2}-2x+4) }{ (x+2) }\]. The x+2's cancel out leaving \[\lim_{x \rightarrow -2} x^{2}-2x+4.\] Now substitute x for -2 and see what happens.

OpenStudy (anonymous):

Oh my goodness, I'm such an idiot. I wasn't expanding x^3. I was keeping it as is. I'm so sorry for all the trouble I put you through. But thank you sooo much! It really is appreciated.

OpenStudy (isaiah.feynman):

I made silly mistake too. Lol.